OpenStudy (anonymous):

solution to the system of equations 2x = 10 - y -2y = 22 - 3x A. (6, -20) B. (6, -2) C. (6, 2) D. (7, -4) HELP MEE

OpenStudy (anonymous):

-y=2x-10 -2y=-3x+22 multiple the first by -2 then add the two equations

OpenStudy (anonymous):

2y=-4x+20 -2y=-3x+22 \[-- -- \] 0=-7x+42 7x=42 x=6

OpenStudy (anonymous):

then y=-2

OpenStudy (anonymous):

Thanks ! im pretty sure i got it down now

OpenStudy (anonymous):

np :)